# On Efficient Variants of Segment Anything Model: A Survey

### v3

@@ 2.2.2 Transformer-alternative Models
State Space Models (SSMs), with near-linear computational complexity relative to sequence length, have emerged as promising alternatives to Transformers. Various methods [130, 131, 132] have been developed to enhance SSMs. One recent advancement, Mamba [133], has made significant strides in processing long sequences. Mamba’s core innovation is the Selective State Spaces, which allow the model to dynamically decide whether to propagate or forget information, enabling more efficient sequence modeling. Additionally, the researchers incorporated a hard-aware algorithm to further accelerate computation on GPUs. Building on this, Vision Mamba [134] applies Mamba to visual representation tasks. Vim introduces Bi-directional SSMs to better capture context within image patch sequences. With its strong performance and efficiency, Vision Mamba is emerging as a potential backbone for vision foundation models.
The Receptance Weighted Key Value (RWKV) model [117], which combines the strengths of both recurrent neural networks (RNNs) and Transformers, achieves linear time complexity as sequence length increases. RWKV is well-suited for handling long sequence processing challenges. To facilitate global information interaction, RWKV replaces the traditional attention mechanism, which has quadratic complexity, with the more efficient WKV Operator and output gating mechanism. These are formulated as follows,
@@
RWKV has also been extended to the vision tasks. The Vision-RWKV (VRWKV) model [135] demonstrates performance comparable to Vision Transformers (ViT) but with greater efficiency. To adapt RWKV from 1D sequences to 2D images, Q-shift tokens is introduced to fuse neighborhood information in four directions. In [136], a RWKV-based variant of SAM has achieved outstanding performance in efficiency, by adopting the efficient backbone mixed of MobileNetV2 blocks [119] and VRWKV blocks.
@@ 2.3.1 Knowledge Distillation
Knowledge distillation (KD) [137] was first introduced as a solution for deploying large, complex neural networks in resource-constrained environments. The core concept is to transfer the knowledge and representation capability from a larger, well-trained model (the teacher model) to a smaller, more efficient model (the student model).
When applying KD to accelerate SAM, the goal is to distill the knowledge from the original, larger SAM and impart it to more efficient SAM-like models. Given SAM’s encoder-decoder architecture, KD can generally be categorized into two approaches: distilling the entire SAM model or distilling the image encoder alone. Most works, such as [138, 43, 113, 44, 122], focus on distilling only the efficient backbone, while retaining the original SAM’s prompt encoder and mask decoder. However, other approaches, like [121, 46], aim to distill the entire model by supervising the outputs of both the encoder and decoder.
@@ 2.3.2 Quantization
where $b,s$ refers to the bit number in low precision and the scaling factor respectively, $Clamp(x,a1,a2)$ clips integer $x$ to $a1/a2$ when $x\leq a1/x\geq a2$. There are two primary types of quantization for neural networks: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ applies quantization to fully trained models using a small calibration dataset to mitigate accuracy loss, while QAT adapts models to low precision during training, requiring the full training dataset. As a result, QAT is generally more costly and inefficient compared to PTQ. Numerous PTQ methods have been developed for CNNs [139, 140, 141] and Transformer-based architectures [142, 143, 144, 145]. In the context of accelerating SAM, works like [44, 47] utilize PTQ techniques to compress SAM with targeted strategies for improved efficiency.
@@ 2.3.3 Pruning
Model pruning reduces a model’s size and complexity by eliminating redundant weights or connections, while aiming to maintain accuracy as much as possible. Pruning methods are typically categorized into two types: structured pruning [146, 147] and unstructured pruning [148, 149].
Structured pruning removes parameters in groups based on specific criteria, targeting substructures such as channels, layers, or blocks in a systematic manner. In contrast, unstructured pruning focuses on individual weights, often resulting in a sparse and fragmented network. However, unstructured pruning may not lead to effective acceleration on general hardware due to the irregularity of the remaining network structure. In [48], structured pruning is applied to lighten SAM, significantly reducing the model’s size while preserving most of SAM’s capabilities by removing large amounts of redundant weights.
@@
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=16em,font=,, where level=2text width=18em,font=,, where level=3text width=18em,font=,, where level=4text width=18em,font=,, where level=5text width=12em,font=,, [Efficient SAM variants, ver [Accelerating SegAny3.1), fill=green!10 [Training from Scratch3.1.1), fill=blue!10 [SAM-different architectures, fill=yellow!10 [FastSAM [41], fill=gray!10] [SqueezeSAM [42], fill=gray!10] ] [SAM-like architectures, fill=yellow!10 [EfficientSAM [109], fill=gray!10] [RMP-SAM [150], fill=gray!10] [SAM 2 [34], fill=gray!10] ] ] [Knowledge Distillation3.1.2), fill=blue!10 [With Light-weight ViT Encoders, fill=yellow!10 [MobileSAM [43], fill=gray!10] [Light HQ-SAM [104], fill=gray!10] [ESAM [121], fill=gray!10] [TinySAM [44], fill=gray!10] ] [With Pure CNN Encoders, fill=yellow!10 [NanoSAM [138], fill=gray!10] [PicoSAM2 [151], fill=gray!10] [RepViT-SAM [45], fill=gray!10] [EdgeSAM [46], fill=gray!10] ] [With Attention-modified Encoders, fill=yellow!10 [EfficientViT-SAM [122], fill=gray!10] [FastSAM3D [128], fill=gray!10] [SAM-Lightening [129], fill=gray!10] [RWKV-SAM [117], fill=gray!10] ] ] [Quantization3.1.3), fill=blue!10 [TinySAM [44], fill=gray!10] [PTQ4SAM [47], fill=gray!10] [PQ-SAM [152], fill=gray!10] ] [Pruning3.1.4), fill=blue!10 [SlimSAM [48], fill=gray!10] [SuperSAM [153], fill=gray!10] ] [Code Refactorization3.1.5), fill=blue!10 [SAMFast [154], fill=gray!10] ] ] [Accelerating SegEvery3.2), fill=green!10 [Panoptic Segmentation Strategy, fill=blue!10 [FastSAM [41], fill=gray!10, tier=number] ] [Efficient Sampling Strategy, fill=blue!10 [MobileSAMV2 [155], fill=gray!10, tier=number] [TinySAM [44], fill=gray!10] [LiteSAM [124], fill=gray!10] [AoP-SAM [156], fill=gray!10] ] ] ]
@@ 2.3.4 Low-Rank Factorization
where $U\in\mathbb{R}^{m\times m},\ V\in\mathbb{R}^{n\times n}$ are orthogonal matrices and $\Sigma\in\mathbb{R}^{m\times n}$ is a diagonal matrix with the none-zero singular values of $A$, while $r\leq\min\{m,n\}$ refers to the rank of $A$. The low-rank factorization has been widely applied in deep neural networks, such as CNNs [157, 158], LSTMs [159, 160], and Transformer-based models [161, 162, 163]. However, to date, no studies have applied this technique to compress and accelerate SAM, presenting a potential future research direction.
Discussion: Sparse matrices derived from low-rank factorization, will take less memory usage and computational cost during forward pass, slimming and accelerating the vanilla model. However, the implementation of some decomposition operations, for instance the SVD, yet being stable and effective in fully connected layers [164], can be computationally expensive especially for large matrices [165], which may hinders its application in SAM. Moreover, determining the best matrix decomposition rank remains to be the crucial challenge for applying factorization approaches. Furthermore, similar to pruning-based methods, decomposing matrices into low rank may result in loss of detailed information, undermining model’s generalization performance. Even though, some recent works applying low-rank factorization in LLMs [166, 167, 168], exploiting the decomposition according to activation distribution or in specific Transformer layers, are proved to be effective, which may provide inspiration to researchers to further implement it on SAM.
@@ 3.1.1 Training from Scratch
Figure 4: The architecture of FastSAM [41]. It takes two stages to achieve segmenting anything: all-instance segmentation and prompt-guided selection. It is worth noting that the outputs of first stage are used directly as the SegEvery results.
@@
FastSAM [41] is one of the first SAM variants that does not rely on SAM’s original Encoder-Decoder architecture. To achieve faster segmentation, it divides the SegAny task into two sub-tasks: all-instance segmentation and prompt-guided selection. Since instance segmentation has been effectively addressed by many CNN-based methods, FastSAM offers improved efficiency compared to the Transformer-based SAM. For instance segmentation, FastSAM employs the YOLOv8-Seg [169] model, using the YOLACT method [170] for enhanced performance. FastSAM can reliably predict objects of interest using points, boxes, or text as prompts. In addition to accelerating the SegAny task, FastSAM also excels in the SegEvery task, as this can be efficiently achieved alongside all-instance segmentation. However, as an early efficient variant of SAM, FastSAM still has some limitations, such as producing low-quality masks for smaller objects and generating masks with less smooth boundaries. Despite these shortcomings, FastSAM marks significant progress by introducing a CNN-based architecture into this domain. The architecture of FastSAM is illustrated in Fig 4.
@@
Building on the successful application of CNNs in SAM, Varadarajan et al. introduced SqueezeSAM [42], which further replaces SAM’s Transformer-based architecture with a U-Net structure [111]. U-Net is composed of an encoder for feature extraction and a decoder for information recovery. SqueezeSAM retains the general U-Net architecture but incorporates two Transformer layers at the lowest scale of the U-Net to strike a balance between speed and accuracy. Additionally, SqueezeSAM features several micro-level optimizations, such as capping the output channels at 256, using BatchNorm [171] instead of LayerNorm [172] for efficiency, and introducing skip connections [112] between the encoder and decoder. A unique challenge for SqueezeSAM lies in handling prompts. Unlike SAM, where prompt tokens are utilized at the decoding stage, SqueezeSAM adopts an early fusion strategy, adding encoded prompts as additional input channels before feeding them into the encoder. The model is trained from scratch using the SA-1B dataset, with data augmentation techniques addressing low-quality data issues. SqueezeSAM is primarily designed for deployment in photography applications, where efficient interactive segmentation is needed. As depicted in Fig. 5, its workflow involves generating an initial mask of the salient object, followed by fine-grained segmentation refined by user clicks.
@@
Figure 5: The architecture of SqueezeSAM [42]. It replaces the Transformer-based encoder-decoder structure with U-Net backbone[111]. Clicks from users and masks of salience objects are fed into SqueezeSAM with the input image to achieve interactive segmentation.
@@
Different from FastSAM and SqueezeSAM, which introduce entirely new structures, EfficientSAM [109] retains SAM’s original architecture but replaces the image encoder. They use ViT-tiny or ViT-small as a lightweight encoder and re-train them from scratch, leveraging the SAM-based Masked Image (SAMI) pretraining strategy. SAMI is adapted from the Masked AutoEncoder (MAE) framework [58], which was initially used to pretrain SAM’s original image encoder. SAMI follows an Encoder-Decoder pipeline: the encoder generates latent feature representations from unmasked patches, while the decoder reconstructs the missing embeddings of masked patches. This process is supervised by reconstruction loss, comparing the embeddings produced by SAM’s ViT-H encoder with those generated by the SAMI pipeline. After pretraining, the lightweight encoder is extracted from the SAMI pipeline and integrated with the rest of SAM’s components, forming EfficientSAM. The final step involves fine-tuning the entire model on the SA-1B dataset for further alignment and refinement. SAMI is a general pretraining approach that can be applied to train any backbone for SAM variants. The overall structure of SAMI and EfficientSAM is illustrated in Fig. 6.
@@
Figure 6: The framework of SAMI and efficientSAM [109]. The target of SAMI is to pretrain a light-weight encoder which holds representation ability close to SAM’s image encoder. And the encoder is further extracted to build up the EfficientSAM.
@@
Concurrent with EfficientSAM, a new variant named RMP-SAM [150] is designed to achieve real-time, all-purpose segmentation, including panoptic segmentation (PS), video instance segmentation (VIS), and interactive segmentation (equivalent to the SegAny task). RMP-SAM retains SAM’s basic Encoder-Decoder architecture but incorporates more efficient components to enhance performance. For the encoder, RMP-SAM combines Feature Pyramid Networks (FPN) [173] with deformable convolutions [174] to extract features from both images and videos, while using a prompt encoder to embed visual prompts. In the decoder, RMP-SAM employs a three-stage pipeline utilizing novel pooling-based dynamic convolutions to refine the mask tokens. The tokens generated in each stage, along with the feature maps from the encoder, serve as inputs. These inputs are first processed by dynamic convolutions [175], followed by refinement using Multi-Head Self Attention (MHSA) and a Feed Forward Network (FFN). After the decoder, two additional prompt adapters are introduced to enhance interaction between the visual prompts and segmentation tokens. The final masks are generated by computing the inner product between the updated tokens and updated prompts. The architecture of RMP-SAM is illustrated in Fig. 7.
@@
Figure 7: The architecture of RMP-SAM[150]. The encoder embeds images and videos to feature maps and encodes visual prompts as queries (tokens). The multi-stage decoder is to refine queries by interacting with feature maps. In the end, the feature maps and refined queries are fed into adapters to generate masks and classes.
@@
Recently, Ravi et al. introduced the Segment Anything Model 2 (SAM 2) [34], an extension of the original SAM, with objective of delivering high-quality, real-time promptable segmentation across both images and videos. In image segmentation tasks, SAM 2 is reported to achieve higher accuracy and a 6× improvement in efficiency compared to the original SAM. This significant advancement is largely attributed to its efficient image encoder, Hiera [59], a hierarchical ViT that has been simplified from MViTv2 [176] by removing redundant components and utilizing the MAE framework for training. Hiera is a streamlined, purely transformer-based architecture that runs faster and delivers better accuracy than traditional ViTs in both image and video tasks.

### v4

@@ 2.2.2 Transformer-alternative Models
State Space Models (SSMs), with near-linear computational complexity relative to sequence length, have emerged as promising alternatives to Transformers. Various methods [gu2021efficiently, gupta2022diagonal, fu2022hungry] have been developed to enhance SSMs. One recent advancement, Mamba [gu2023mamba], has made significant strides in processing long sequences. Mamba’s core innovation is the Selective State Spaces, which allow the model to dynamically decide whether to propagate or forget information, enabling more efficient sequence modeling. Additionally, the researchers incorporated a hard-aware algorithm to further accelerate computation on GPUs. Building on this, Vision Mamba [zhu2024vision] applies Mamba to visual representation tasks. Vim introduces Bi-directional SSMs to better capture context within image patch sequences. With its strong performance and efficiency, Vision Mamba is emerging as a potential backbone for vision foundation models.
The Receptance Weighted Key Value (RWKV) model [peng2023rwkv], which combines the strengths of both recurrent neural networks (RNNs) and Transformers, achieves linear time complexity as sequence length increases. RWKV is well-suited for handling long sequence processing challenges. To facilitate global information interaction, RWKV replaces the traditional attention mechanism, which has quadratic complexity, with the more efficient WKV Operator and output gating mechanism. These are formulated as follows,
@@
RWKV has also been extended to the vision tasks. The Vision-RWKV (VRWKV) model [duan2024vision] demonstrates performance comparable to Vision Transformers (ViT) but with greater efficiency. To adapt RWKV from 1D sequences to 2D images, Q-shift tokens is introduced to fuse neighborhood information in four directions. In [yuan2024mamba], a RWKV-based variant of SAM has achieved outstanding performance in efficiency, by adopting the efficient backbone mixed of MobileNetV2 blocks [sandler2018mobilenetv2] and VRWKV blocks.
@@ 2.3.1 Knowledge Distillation
Knowledge distillation (KD) [hinton2015distilling] was first introduced as a solution for deploying large, complex neural networks in resource-constrained environments. The core concept is to transfer the knowledge and representation capability from a larger, well-trained model (the teacher model) to a smaller, more efficient model (the student model).
When applying KD to accelerate SAM, the goal is to distill the knowledge from the original, larger SAM and impart it to more efficient SAM-like models. Given SAM’s encoder-decoder architecture, KD can generally be categorized into two approaches: distilling the entire SAM model or distilling the image encoder alone. Most works, such as [nanosam, zhang2023faster, wang2023repvit, shu2023tinysam, zhang2024efficientvit], focus on distilling only the efficient backbone, while retaining the original SAM’s prompt encoder and mask decoder. However, other approaches, like [Zhao2023ESAM, zhou2023edgesam], aim to distill the entire model by supervising the outputs of both the encoder and decoder.
@@ 2.3.2 Quantization
where $b,s$ refers to the bit number in low precision and the scaling factor respectively, $Clamp(x,a1,a2)$ clips integer $x$ to $a1/a2$ when $x\leq a1/x\geq a2$. There are two primary types of quantization for neural networks: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ applies quantization to fully trained models using a small calibration dataset to mitigate accuracy loss, while QAT adapts models to low precision during training, requiring the full training dataset. As a result, QAT is generally more costly and inefficient compared to PTQ. Numerous PTQ methods have been developed for CNNs [banner2019post, nahshan2021loss, wang2024aqa] and Transformer-based architectures [frantar2022gptq, yao2022zeroquant, liu2021post, yuan2022ptq4vit]. In the context of accelerating SAM, works like [shu2023tinysam, lv2024ptq4sam] utilize PTQ techniques to compress SAM with targeted strategies for improved efficiency.
@@ 2.3.3 Pruning
Model pruning reduces a model’s size and complexity by eliminating redundant weights or connections, while aiming to maintain accuracy as much as possible. Pruning methods are typically categorized into two types: structured pruning [anwar2017structured, wang2019structured] and unstructured pruning [laurent2020revisiting, chen2020tight].
Structured pruning removes parameters in groups based on specific criteria, targeting substructures such as channels, layers, or blocks in a systematic manner. In contrast, unstructured pruning focuses on individual weights, often resulting in a sparse and fragmented network. However, unstructured pruning may not lead to effective acceleration on general hardware due to the irregularity of the remaining network structure. In [chen20230], structured pruning is applied to lighten SAM, significantly reducing the model’s size while preserving most of SAM’s capabilities by removing large amounts of redundant weights.
@@
Efficient SAM variants
Accelerating SegAny3.1) Training from Scratch3.1.1) SAM-different architectures FastSAM [zhao2023fast] SqueezeSAM [varadarajan2023squeezesam] SAM-like architectures EfficientSAM [xiong2024efficientsam] RMP-SAM [xu2024rap] SAM 2 [ravi2024sam] Knowledge Distillation3.1.2) With Light-weight ViT Encoders MobileSAM [zhang2023faster] Light HQ-SAM [ke2024segment] ESAM [Zhao2023ESAM] TinySAM [shu2023tinysam] With Pure CNN Encoders NanoSAM [nanosam] PicoSAM2 [bonazzi2025picosam2] RepViT-SAM [wang2024repvit] EdgeSAM [zhou2023edgesam] With Attention-modified Encoders EfficientViT-SAM [zhang2024efficientvit] FastSAM3D [shen2024fastsam3d] SAM-Lightening [songa2024sam] RWKV-SAM [peng2023rwkv] Quantization3.1.3) TinySAM [shu2023tinysam] PTQ4SAM [lv2024ptq4sam] PQ-SAM [liu2024pq] Pruning3.1.4) SlimSAM [chen20230] SuperSAM [abebe2025supersam] Code Refactorization3.1.5) SAMFast [SAMfast] Accelerating SegEvery3.2) Panoptic Segmentation Strategy FastSAM [zhao2023fast] Efficient Sampling Strategy MobileSAMV2 [zhang2023mobilesamv2] TinySAM [shu2023tinysam] LiteSAM [fu2024lite] AoP-SAM [chenaop]
@@ 2.3.4 Low-Rank Factorization
where $U\in\mathbb{R}^{m\times m},\ V\in\mathbb{R}^{n\times n}$ are orthogonal matrices and $\Sigma\in\mathbb{R}^{m\times n}$ is a diagonal matrix with the none-zero singular values of $A$, while $r\leq\min\{m,n\}$ refers to the rank of $A$. The low-rank factorization has been widely applied in deep neural networks, such as CNNs [tai2015convolutional, yu2017compressing], LSTMs [grachev2017neural, winata2019effectiveness], and Transformer-based models [tahaei2021kroneckerbert, edalati2021kronecker, hajimolahoseini2022strategies]. However, to date, no studies have applied this technique to compress and accelerate SAM, presenting a potential future research direction.
Discussion: Sparse matrices derived from low-rank factorization, will take less memory usage and computational cost during forward pass, slimming and accelerating the vanilla model. However, the implementation of some decomposition operations, for instance the SVD, yet being stable and effective in fully connected layers [xue2013restructuring], can be computationally expensive especially for large matrices [swaminathan2020sparse], which may hinders its application in SAM. Moreover, determining the best matrix decomposition rank remains to be the crucial challenge for applying factorization approaches. Furthermore, similar to pruning-based methods, decomposing matrices into low rank may result in loss of detailed information, undermining model’s generalization performance. Even though, some recent works applying low-rank factorization in LLMs [yuan2023asvd, sharma2023truth, zhu2024survey], exploiting the decomposition according to activation distribution or in specific Transformer layers, are proved to be effective, which may provide inspiration to researchers to further implement it on SAM.
@@ 3.1.1 Training from Scratch
Figure 4: The architecture of FastSAM [zhao2023fast]. It takes two stages to achieve segmenting anything: all-instance segmentation and prompt-guided selection. It is worth noting that the outputs of first stage are used directly as the SegEvery results.
@@
FastSAM [zhao2023fast] is one of the first SAM variants that does not rely on SAM’s original Encoder-Decoder architecture. To achieve faster segmentation, it divides the SegAny task into two sub-tasks: all-instance segmentation and prompt-guided selection. Since instance segmentation has been effectively addressed by many CNN-based methods, FastSAM offers improved efficiency compared to the Transformer-based SAM. For instance segmentation, FastSAM employs the YOLOv8-Seg [yolov8] model, using the YOLACT method [bolya2019yolact] for enhanced performance. FastSAM can reliably predict objects of interest using points, boxes, or text as prompts. In addition to accelerating the SegAny task, FastSAM also excels in the SegEvery task, as this can be efficiently achieved alongside all-instance segmentation. However, as an early efficient variant of SAM, FastSAM still has some limitations, such as producing low-quality masks for smaller objects and generating masks with less smooth boundaries. Despite these shortcomings, FastSAM marks significant progress by introducing a CNN-based architecture into this domain. The architecture of FastSAM is illustrated in Fig 4.
@@
Building on the successful application of CNNs in SAM, Varadarajan et al. introduced SqueezeSAM [varadarajan2023squeezesam], which further replaces SAM’s Transformer-based architecture with a U-Net structure [ronneberger2015u]. U-Net is composed of an encoder for feature extraction and a decoder for information recovery. SqueezeSAM retains the general U-Net architecture but incorporates two Transformer layers at the lowest scale of the U-Net to strike a balance between speed and accuracy. Additionally, SqueezeSAM features several micro-level optimizations, such as capping the output channels at 256, using BatchNorm [ioffe2015batch] instead of LayerNorm [ba2016layer] for efficiency, and introducing skip connections [he2016deep] between the encoder and decoder. A unique challenge for SqueezeSAM lies in handling prompts. Unlike SAM, where prompt tokens are utilized at the decoding stage, SqueezeSAM adopts an early fusion strategy, adding encoded prompts as additional input channels before feeding them into the encoder. The model is trained from scratch using the SA-1B dataset, with data augmentation techniques addressing low-quality data issues. SqueezeSAM is primarily designed for deployment in photography applications, where efficient interactive segmentation is needed. As depicted in Fig. 5, its workflow involves generating an initial mask of the salient object, followed by fine-grained segmentation refined by user clicks.
@@
Figure 5: The architecture of SqueezeSAM [varadarajan2023squeezesam]. It replaces the Transformer-based encoder-decoder structure with U-Net backbone[ronneberger2015u]. Clicks from users and masks of salience objects are fed into SqueezeSAM with the input image to achieve interactive segmentation.
@@
Different from FastSAM and SqueezeSAM, which introduce entirely new structures, EfficientSAM [xiong2024efficientsam] retains SAM’s original architecture but replaces the image encoder. They use ViT-tiny or ViT-small as a lightweight encoder and re-train them from scratch, leveraging the SAM-based Masked Image (SAMI) pretraining strategy. SAMI is adapted from the Masked AutoEncoder (MAE) framework [he2022masked], which was initially used to pretrain SAM’s original image encoder. SAMI follows an Encoder-Decoder pipeline: the encoder generates latent feature representations from unmasked patches, while the decoder reconstructs the missing embeddings of masked patches. This process is supervised by reconstruction loss, comparing the embeddings produced by SAM’s ViT-H encoder with those generated by the SAMI pipeline. After pretraining, the lightweight encoder is extracted from the SAMI pipeline and integrated with the rest of SAM’s components, forming EfficientSAM. The final step involves fine-tuning the entire model on the SA-1B dataset for further alignment and refinement. SAMI is a general pretraining approach that can be applied to train any backbone for SAM variants. The overall structure of SAMI and EfficientSAM is illustrated in Fig. 6.
@@
Figure 6: The framework of SAMI and efficientSAM [xiong2024efficientsam]. The target of SAMI is to pretrain a light-weight encoder which holds representation ability close to SAM’s image encoder. And the encoder is further extracted to build up the EfficientSAM.
@@
Concurrent with EfficientSAM, a new variant named RMP-SAM [xu2024rap] is designed to achieve real-time, all-purpose segmentation, including panoptic segmentation (PS), video instance segmentation (VIS), and interactive segmentation (equivalent to the SegAny task). RMP-SAM retains SAM’s basic Encoder-Decoder architecture but incorporates more efficient components to enhance performance. For the encoder, RMP-SAM combines Feature Pyramid Networks (FPN) [lin2017feature] with deformable convolutions [dai2017deformable] to extract features from both images and videos, while using a prompt encoder to embed visual prompts. In the decoder, RMP-SAM employs a three-stage pipeline utilizing novel pooling-based dynamic convolutions to refine the mask tokens. The tokens generated in each stage, along with the feature maps from the encoder, serve as inputs. These inputs are first processed by dynamic convolutions [chen2020dynamic], followed by refinement using Multi-Head Self Attention (MHSA) and a Feed Forward Network (FFN). After the decoder, two additional prompt adapters are introduced to enhance interaction between the visual prompts and segmentation tokens. The final masks are generated by computing the inner product between the updated tokens and updated prompts. The architecture of RMP-SAM is illustrated in Fig. 7.
@@
Figure 7: The architecture of RMP-SAM[xu2024rap]. The encoder embeds images and videos to feature maps and encodes visual prompts as queries (tokens). The multi-stage decoder is to refine queries by interacting with feature maps. In the end, the feature maps and refined queries are fed into adapters to generate masks and classes.
@@
Recently, Ravi et al. introduced the Segment Anything Model 2 (SAM 2) [ravi2024sam], an extension of the original SAM, with objective of delivering high-quality, real-time promptable segmentation across both images and videos. In image segmentation tasks, SAM 2 is reported to achieve higher accuracy and a 6× improvement in efficiency compared to the original SAM. This significant advancement is largely attributed to its efficient image encoder, Hiera [ryali2023hiera], a hierarchical ViT that has been simplified from MViTv2 [li2022mvitv2] by removing redundant components and utilizing the MAE framework for training. Hiera is a streamlined, purely transformer-based architecture that runs faster and delivers better accuracy than traditional ViTs in both image and video tasks.